OpenStudy (aliaaalihassan):

Find the generalized form of the integral: (X^n)/(1+x^2)

OpenStudy (aliaaalihassan):

@wio

OpenStudy (aliaaalihassan):

@shrutipande9

OpenStudy (aliaaalihassan):

for any positive integer n,

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

use trig substitution as x= tan theta. so \[\int\limits_{?}^{?}\frac{ x^n }{ 1+x^2 }dx = \int\limits_{?}^{?} \tan^n \theta d \theta=\int\limits_{?}^{?}(1+\sec^2)\tan ^{n-2}\theta d \theta\] I you can continue to get the final form.

OpenStudy (michele_laino):

Please in order to get your answer, it is necessary to evaluate your integral for n=0,1 and then we have to consider the cases n odd and n even. I got your solution, nevertheless I can not post it since the Code-of Conduct

OpenStudy (anonymous):

excuse me @Michele_Laino what is Code-of Conduct?

OpenStudy (michele_laino):

please see the link below: http://openstudy.com/code-of-conduct @Catch.me

OpenStudy (aliaaalihassan):

@michele_laino can you solve it for me?

OpenStudy (aliaaalihassan):

\[\int\limits_{0}^{1}\frac{ 1 }{1+x^2} dx =\tan^{-1} x \]

OpenStudy (aliaaalihassan):

\[\int\limits_{0}^{1}\frac{ x }{ 1+x^2 } dx= \frac{ 1 }{ 2 }\ln (1+x^2)\]

OpenStudy (aliaaalihassan):

\[\int\limits_{0}^{1}\frac{ x^2 }{ 1+x^2} dx = x-\tan^{-1} x\]

OpenStudy (aliaaalihassan):

\[\int\limits_{0}^{1} \frac{ x^3 }{ 1+x^2 } dx = 1/2x^2 -1/2\ln(x^2+1)\]

OpenStudy (anonymous):

For \(n\ge3\), you have the recursive formula \[I_n=\int\frac{x^n}{1+x^2}\,dx=\int\left(x^{n-2}-\frac{x^{n-2}}{1+x^2}\right)\,dx=\int x^{n-2}\,dx-I_{n-2}\] Judging by the alternating pattern between \(\arctan\) and \(\ln\) terms above, I would expect a different result for even and odd \(n\). Try breaking up into two cases, with \(k=1,2,3,...\) (such that \(2k+1=3\) and \(2k=4\), and so on): \[I_{2k}=\int\frac{x^{2k}}{1+x^2}\,dx\quad\text{and}\quad I_{2k+1}=\int\frac{x^{2k+1}}{1+x^2}\,dx\] For the latter, a substitution of \(u=1+x^2\) gives \(\dfrac{du}{2}=x\,dx\), and so \[I_{2k+1}=\frac{1}{2}\int \frac{(u-1)^k}{u}\,du\] This can be expanded using the binomial theorem and integrated term by term. For the even integral, setting \(x=\tan t\) gives \(dx=\sec^2t\,dt\), hence \[\begin{align*}I_{2k}&=\int\frac{\tan^{2k}t\sec^2t}{1+\tan^2t}\,dt\\\\&=\int\tan^{2k}t\,dt\\\\&=\int(\tan^{2k-2}t\sec^2t-\tan^{2k-2}t)\,dt\\\\ &=\frac{\tan^{2k-1}t}{2k-1}-\int\tan^{2k-2}t\,dt\\\\ I_{2k}&=\frac{\tan^{2k-1}t}{2k-1}-I_{2k-2}\end{align*}\]